OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (anonymous):

exp(5) = x + 1 x = exp(5) -1 = 148.4 -1 = 147.4 check ln(147.4 + 1) = 5 yes

OpenStudy (anonymous):

x=147.41 to nearest hundredth